Consider a monopolist facing the daily demand curve displayed below. Use the line tool to draw the marginal revenue curve.

To refer to the graphing tutorial for this question type, please click here


Price 16 15 14 13 12 11 10 9 7 6 5 4 3 2 D 10600 1100 1200 1300 1400 1500 1600 Quantity

Part 2   (1 point)

See Hint

Suppose that the monopolist's cost function is given by ?(?)=4?c(y)=4y, where ?y stands for the amount of output produced daily. The firm's marginal cost equals $ _____   .

Part 3   (1 point)

See Hint

If the monopolist is a profit maximizer, the monopolist will produce and sell _____   units per day. (Round your answer to two decimals if necessary.)

Part 4   (1 point)

See Hint

If this is a profit-maximizing monopolist, the price per unit of the product will be $ _____ . (Round your answer to the nearest integer.)
Answer #1

The Demand Curve can be drawn as follows

(1400 - y) / 1400 = (0 - P) / (0 - 12)

or 1 - y/1400 = P / 1200

or 12 - 3/350*y = P

The marginal Revenue curve is

MR = 12 - 3/175y

The graph can be drawn as follows

Monopolist Chart 14 12 10 8 Price 6 4 2 0 0 100 200 300 400 500 600 700 800 900 1000 1100 1200 1300 1400 Quantity Demand Curv

Part 2:

Cost function is given by c(y) = 4y where y is the output

Marginal cost = d/dy (c(y)) = d/dy (4y) = 4

Suppose that the monopolist's cost function is given by ?(?)=4?c(y)=4y, where ?y stands for the amount of output produced daily. The firm's marginal cost equals $ 4

Part 3:

At profit maximization, MR = MC

or 12 - 3/175y = 4

or 3/175y = 8

or y = 466.67

So the monopolist will sell 466.67 units per day

If the monopolist is a profit maximizer, the monopolist will produce and sell 466.67 units per day.

Part 4:

The price of each unit at profit maximization is given as

P = 12 - 3/350*y = 12 - 3*466.67 / 350 = 8

If this is a profit-maximizing monopolist, the price per unit of the product will be $ 8

If you found this helpful, please rate it so that I can have higher earnings at no extra cost to you. This will motivate me to write more.